The first year student in the school of health sciences was asked to prepare 250cm3 of 0.1mol/dm3 sodium carbonate solution. The student weighed 2.65g of sodium carbonate decahydrate, (Na2CO3.10H2O), instead of anhydrous sodium carbonate (Na2CO3). Calculate the concentration of the solution prepared by the student in mol/dm3
1
Expert's answer
2019-09-12T03:33:01-0400

MNa2CO3x10H2O = 286 g/mol

CNa2CO3= = m/(MNa2CO3x10H2O* V) = 2.65/(286*0.25) = 0.037 mol/dm^3